# Memento: Towards Proactive Visualization of Everyday Memories with Personal Wearable AR Assistant

### v1

@@ 2.2.3 Personal Memory Generation
The referent is at the heart of the memory that bridges AR visualization with the physical realm (Fig. 1). Upon initial query interaction with a physical referent via Personal assistant, the query and its metadata are recorded in the form of a memory. It holds the moment of verbal query, response, source of response, and its context of the spatial activity. We term this, Referent-anchored Spatiotemporal Activity Memory (RSAM).
To generate an RSAM, we first identify the referent in question, and retrieve the most recent Spatial Activity Profile. Then, we bring the user’s verbal query, its response, source of response, all together, and form a memory. We store a new RSAM in a data structure that combines R-tree [rtree] with a graph-based hierarchical $K$-approximate nearest neighbor search algorithm known as HNSW [malkov2018efficient]. This hybrid structure enables us fast insertion and retrieval of RSAM, suitable for interactive AR assistance system. The R-tree provides coarse partitioning of data based on latitude, longitude, timestamp keys, while the fast fetching of vector-represented spatial activity context, enables the capturing of the semantics of user’s scene and activity via HNSW.
@@ 2.2.4 Context-aware Memory Retrieval
Recalling of a memory begins from accessing the most recent Spatial Activity Profile. It contains the user’s current geographical location, scene information, activity, and timestamp. By filtering the existing RSAMs by geographic coordinates and time, first, and hierarchically retrieving vectorized spatial activity, Memento enables fast top-$K$ context similarity retrieval in a deterministic way.
In an effort to reduce false positive memory suggestions upon the retrieval of RSAMs, we equally weigh the significance of each dimension in Spatial Activity Profile. We sequentially verify the semantic alignment between each dimension, and user’s current context. The proximity of spatiotemporal context is done as the first step, then, the existence of the RSAM referents in user’s view, and fine-grained spatial activity semantics matching. Only when all dimensions are satisfied, an RSAM is suggested to the user.
@@ 2.2.5 Proactive Situated Information Display
Referent detection. Once the $K$-similar RSAMs–the past memories that align with the user’s current spatiotemporal and activity context–return, Proactive visualizer localizes the referents where the memories were situated onto. This situation of AR content is done by localizing the referents specified in RSAMs via open-vocabulary detector. If present the current view, the 2D center points are transmitted to the user’s HMD to compute the 3D anchor points using the depth map acquired from the external camera of the HMD, at the time of the image capture.
Content visualization. A memory is AR anchored onto the referent an RSAM data specifies. As shown in Fig. 3-A, users are given the option to view the system reasoning on Spatial Activity contextual alignment. The Space, Time, and Activity of the matched information in a human-readable format is shown for explainability. ‘Remove’ option allows permanent dismissal of unwanted (e.g., miss-matched context) or a single-time query that does not require repeating. Memento also proactively self-update the data of the contents providing newly available information on each repeat cycle. This makes Memento, beyond a simple immersive reminder, but a fresh information fetching agent.
Content interaction. To interact intuitively with the memories, we employ “head-gaze-and-pinch” and the design pattern of “Head-gaze and dwell” [microsoftheadgazedwell]. Users can naturally look towards (head-gaze) the memory of interest, and select it with the pinching gesture. It not only provides embodied way of interacting with data, but also, a double safe measure to confirm the desired memory of interest before initiating the raycast selection.
@@ 2.2.6 User Preference and Proactivity Adjustment Interface
Fully autonomous information suggestion systems risk diminishing the user’s sense of ownership and control [liu2023visual]. As discussed in Section 2.2.5, Memento presents a user interface that indicates the spatial activity of the user’s situated memory (space, time, activity), and the option to remove the memory (Fig. 3-A). Furthermore, the level of proactivity of AR visualizations can vary depending on personal preferences and tasks [meurisch2020exploring], thus, we provide an interface to adjust the memory recall sensitivity (Fig. 3-B) that controls the frequency of the recalls (How often the system recalls), in days.
@@ 3 Example Use Cases and Applications
Memento leverages recurring patterns of daily activities to proactively surface relevant information without requiring repeated explicit queries. Rather than prompting the user to ask again, the system infers when past needs are likely to reappear based on recurring spatiotemporal routines and resurfaces the most contextually relevant memories through situated visualization on everyday objects. Below are three representative patterns of daily use, observed from our user study (denoted A1 to A3 across the sections and in Fig. 4).
@@
Staying up-to-date in information-rich contexts (A2). For activities involving regular information monitoring, such as research and entertainment, users often revisit similar referents and spaces with similar intentions (e.g., sitting at a workstation to look up new papers, or taking a break in a coffee corner while catching up on sports news). Recognizing these consistent contextual cues, the system retrieves previously asked queries (e.g., “latest research papers”) and presents fresh responses situated on familiar objects in those spaces. This pairing of personal querying history with recurring usage locations turns routine contexts into lightweight information provider.
@@ 4 Preliminary Evaluation
To assess the practicality and potential of proactive, context-aware AR assistance, we conducted a two-part preliminary evaluation: (1) a technical analysis of Memento the capabilities and performance, and (2) a “take-home” diary-based user study, examining user acceptance, user experience, real-world usage patterns, and how the technology integrates into everyday routines.
@@ 4.1 Technical Evaluation
Identifying users’ action contexts, and delivering timely and accurate situated AR content, are crucial for the practical adoption of proactive context-aware systems. In this section, we assess context-aware capabilities of Memento, and report them.
Spatial activity context matching accuracy. We evaluate the accuracy of the Spatial Activity Recognizer in identifying spatial and activity contexts from egocentric views using text-based scene representations. Each image is converted into a textual description with GPT-4o, and CLIP is used to generate embeddings directly from these textual summaries. To assess classification performance, we compute centroid embeddings for each spatial and activity class category using a sampled dataset of 1,000 egocentric images drawn from Ego4D. For each test sample, its embedding is compared against class centroids using cosine similarity, retrieving the closest match as the predicted label.
The recognizer achieves an F1 score of 0.82 for Scene context classification and 0.82 for Activity context classification (see Table 1). We find that this tendency continues until images contains motion blurs or low lighting conditions, where the accuracy drops to an F1 score of 0.57 and 0.30, respectively. These results indicate that text-based embeddings can provide semantic distinction to identify recurring spatial activities for proactive resurfacing, even without fine-tuning or visual modality fusion.
Interaction referent recognition accuracy. We found that the open-vocabulary object detector, YOLO-World-XL, occasionally struggles to robustly localize a referent the user is interacting with. While it outperforms DetCLIP-T in speed (52 vs 2FPS) with similar accuracy (34.4 vs 35.4 AP), false positives led to faulty retrieval of RSAM and inconsistent situation of proactive AR visualizations, impacting the user experience.
@@
Table 1: Performance comparison of CLIP variants–CNN-based (ResNet-101) and Transformer-based (ViT-B/32, ViT-L/14) architectures, for for spatial Scene and Activity recognition tasks.
@@ 4.2 User Evaluation and Feedback
We collected data from 25 sessions conducted over 2-4 day period, resulting in a total of $\approx$435 aggregated minutes, and 196 memories (total number of query-generated memories across sessions; including user-dismissed memories) across all participants. Although each session lasted for 17.4 minutes in average, being relatively short in duration, participants were exposed to 135 memories via proactive recalls (68.9% of total memories).
@@
Participants and procedure. We recruited 9 (P1-P9) self-volunteered participants (7 male, 2 female), aged 25-36 ($\mu$=29.1, $\sigma$=3.7), from Computer Science, Mathematics, and Physics background. Prior AR familiarity was broadly distributed on a 5-point scale (1:Not familiar, 5:Familiar), ($\mu$=2.8, $\sigma$=1.3), and gesture-based interaction experience showed similar spread ($\mu$=3.1, $\sigma$=1.6). All participants reported regular daily use of personal computing devices. The identity of the participants was anonymized, and each provided informed consent prior to the participation. This study was conducted under the IRB compliance of Stony Brook University ($1173920$). The participants were instructed to complete a brief pre-study questionnaire and a 30-minute introduction to Memento. They were told to use Memento naturally over 2-4 days in their daily indoor environments, with each usage counted as a ‘session.’ After each session, participants submitted short diary entries describing context, perceived usefulness, and ease of integration, including a 5-point Likert scale and open-ended feedback. System logs of spatial context, interactions, and resurfaced memories were collected in parallel. After its multi-day use, participants completed a post-study survey comparing proactive and reactive support. Semi-structured interviews followed to explore how proactive cues affected routines and what limitations they encountered. To ensure consistency across participants, proactive resurfacing sensitivity was fixed at one resurfacing per 24 hours.
@@
Repeated daily usages. Participants praised Memento for its proactive task reminding ability–“I liked that I could place reminders around the house. Next to my shoes, I’ll know I need to go out for a run”(P3). However, one user reported referent misalignment: “It mistook a square grocery bag to be a luggage and suggested me travel time”(P3). Post-study interview revealed that the user closed up to the referent, enforcing our recognition model to infer with minimal visual context. In another case, a spatial activity was over-generalized: “It said I’m gazing at an object. It’s too generic”(P2). We discuss the candidate solutions to the stated issues in Section 5.
@@
Acceptability in daily routine. Users were favorable towards Memento becoming part of their daily routine: “Each time I use it, I get more comfortable with the interaction style. It feels like a solid foundation”(P5). However, the requests to extend Memento, to a lightweight glasses-based AR were suggested across participants (P1, P3, P4, P6, P8); “It’ll be nice to have a VR app integrated into lifestyle through glasses that pops up information quick”(P6). A non-tech user raised : “It is cool, but I wear a huge glasses, so I can’t wear it for long”(P3). Furthermore, even with a multi-day AR experience, users expressed the need for a longer duration, to observe any changes in their routines: “I cannot say it changed my routines yet. It clearly helped me, but I’d need to try it on for longer. I also might feel differently if it is a glasses I can wear casually, like the Ray-ban”(P1); “Very promising system. Looking forward to seeing how it adapts with more use”(P5). We discuss on the extension of Memento in Section 5.
@@
Proactive vs Reactive assistance. Proactive context-aware AR of Memento were evaluated positively: “I check weather and traffic every morning, this gave me the answers before I checked”(P7); “It reduces a step in your daily patterns”(P6); “Great for reminders. I usually use Google Calendar, but many times my phone is on silent or put away, I’m unable to follow through. Proactivity definitely helps more and put bunch of To-Dos around my house”(P3). Another participant compared Memento to a traditional reactive assistant: “I use Google Nest Mini to voice query and get responses, but Memento recall of previous queries is very useful”(P1), and praised on its simplicity “Biggest pros is that I don’t need to think about setting up the conditions, just automatic and simple”(P6).
5 Future Works and Discussion
@@
Mixed-initiative context control and activity granularity. Accurate resurfacing depends on inferring when a resurfaced memory is contextually appropriate. An AR content that is mal-triggered may negate the usefulness and the utility of proactive visualization. While participants generally appreciated autonomous triggers, they wanted occasional corrective control: “I sometimes wanted more control”(P5). This suggests that proactive AR should adopt mixed-initiative mechanisms that remain simple, yet adjustable. Prior work on user-assistant collaboration and conditional triggers, such as IFTTT (If-This-Then-That) [ifttt, shakeri2023user]-like structures, may provide contexts for lightweight refinements. Improving recognition accuracy may also benefit from hierarchical activity representations as suggested by ongoing efforts in structured activity inference [edge2024local, userllm, xie2024embodied, yang2025egolife], mitigating over-generalization and help maintain alignment between resurfacing timing and user intent.
Continuity of context across realities. Participants responded positively when resurfaced information appeared, “where it felt natural,” especially when memories were tied to objects already associated with an activity. While the current panel-based overlays provide invariance to layout changes, they lack deeper geometric coupling with the environment. Closer integration between AR content and spatial referents could strengthen relevance and improve interpretability and explainability, building on design considerations for spatially anchored AR interfaces [lee2023design]. Future work should explore referent-aware embedding that balances spatial robustness with meaningful anchoring.
Need for longitudinal routine pattern evaluation. Our diary-based deployment captured short-term routine patterns but could not reveal how resurfacing should adapt as habits evolve. Participants occasionally treated the system as a temporary “study task,” leading to usage that did not fully reflect everyday rhythm. Longer-term field studies would allow users to internalize proactive AR as part of daily life, similar to long-term wearable research [bressa2022data]. Understanding when proactive reminders should scale back, pause, or revive requires extended observation of routine drift and habit change over weeks or months, rather than days.
@@ 6 Conclusion
To explore the usability of proactive, context-aware assistance on a wearable AR device, we designed a groundwork system, Memento, that captures, stores users’ verbal interactions as Referent-anchored Spatiotemporal Activity Memories, and proactively resurfaces them based on recurring daily contexts. We introduced a hybrid data retrieval structure combining spatial indexing and semantic vector search, and implemented a simplified context-aware system by integrating multimodal sensing, LLM-based reasoning, and open-vocabulary referent detection. Through real-world use cases, we investigated the potential of everyday use of proactive AR assistant, and discuss takeaways and future works.

### v2

@@ 2.2.3 Personal Memory Generation
The referent is at the heart of the memory that bridges AR visualization with the physical realm (Fig. 1). Upon initial query interaction with a physical referent via Personal assistant, the query and its metadata are recorded in the form of a memory. It holds the moment of the verbal query, the response, the source of the response, and the context of the spatial activity. We term this, Referent-anchored Spatiotemporal Activity Memory (RSAM).
To generate an RSAM, we first identify the referent in question, and retrieve the most recent Spatial Activity Profile. Then, we bring the user’s verbal query, its response, and the response source together, and form a memory. We store a new RSAM in a data structure that combines R-tree [rtree] with a graph-based hierarchical $K$-approximate nearest neighbor search algorithm known as HNSW [malkov2018efficient]. This hybrid structure enables fast insertion and retrieval of RSAM, suitable for an interactive AR assistance system. The R-tree provides coarse partitioning of data based on latitude, longitude, and timestamp keys, while fast retrieval of the vector-represented spatial activity context captures the semantics of the user’s scene and activity via HNSW.
@@ 2.2.4 Context-aware Memory Retrieval
Memory recall begins by accessing the most recent Spatial Activity Profile. It contains the user’s current geographical location, scene information, activity, and timestamp. By filtering the existing RSAMs by geographic coordinates and time first, and hierarchically retrieving vectorized spatial activity, Memento enables fast top-$K$ context similarity retrieval in a deterministic way.
In an effort to reduce false positive memory suggestions upon the retrieval of RSAMs, we equally weigh the significance of each dimension in the Spatial Activity Profile. We sequentially verify the semantic alignment between each dimension, and the user’s current context. The proximity of spatiotemporal context is assessed as the first step, then we verify the existence of the RSAM referents in the user’s view, and finally fine-grained spatial activity semantics matching. Only when all dimensions are satisfied, an RSAM is suggested to the user.
@@ 2.2.5 Proactive Situated Information Display
Referent detection. Once the $K$-similar RSAMs–the past memories that align with the user’s current spatiotemporal and activity context–return, the Proactive visualizer localizes the referents where the memories were situated previously. This situating of AR content is done by localizing the referents specified in RSAMs via an open-vocabulary detector. If present in the current view, the 2D center points are transmitted to the user’s HMD to compute the 3D anchor points using the depth map acquired from the external camera of the HMD, at the time of the image capture.
Content visualization. A memory is AR-anchored onto the referent specified by an RSAM. As shown in Fig. 3-A, users are given the option to view the system reasoning on Spatial Activity contextual alignment. The Space, Time, and Activity of the matched information in a human-readable format is shown for explainability. The ‘Remove’ option allows permanent dismissal of unwanted (e.g., mismatched) memories or one-time queries that do not require repeating. Memento also proactively self-update the data of the contents providing newly available information on each repeat cycle. This makes Memento, beyond a simple immersive reminder, a fresh information-fetching agent.
Content interaction. To interact intuitively with the memories, we employ “head-gaze-and-pinch” and the design pattern of “Head-gaze and dwell” [microsoftheadgazedwell]. Users can naturally look towards (head-gaze) the memory of interest, and select it with the pinch gesture. It not only provides an embodied way of interacting with data, but also serves as a two-step safeguard to confirm the desired memory of interest before initiating the raycast selection.
@@ 2.2.6 User Preference and Proactivity Adjustment Interface
Fully autonomous information suggestion systems risk diminishing the user’s sense of ownership and control [liu2023visual]. As discussed in Section 2.2.5, Memento presents a user interface that indicates the spatial activity of the user’s situated memory (space, time, activity), and the option to remove the memory (Fig. 3-A). Furthermore, the level of proactivity of AR visualizations can vary depending on personal preferences and tasks [meurisch2020exploring], thus, we provide an interface to adjust the memory recall sensitivity (Fig. 3-B) that controls the frequency of recall (How often the system recalls), in days.
@@ 3 Example Use Cases and Applications
Memento leverages recurring patterns of daily activities to proactively surface relevant information without requiring repeated explicit queries. Rather than prompting the user to ask again, the system infers when past needs are likely to reappear based on recurring spatiotemporal routines and resurfaces the most contextually relevant memories through situated visualization on everyday objects. Below are three representative patterns of daily use, observed in our user study (denoted A1 to A3 across the sections and in Fig. 4).
@@
Staying up-to-date in information-rich contexts (A2). For activities involving regular information monitoring, such as research and entertainment, users often revisit similar referents and spaces with similar intentions (e.g., sitting at a workstation to look up new papers, or taking a break in a coffee corner while catching up on sports news). Recognizing these consistent contextual cues, the system retrieves previously asked queries (e.g., “latest research papers”) and presents fresh responses situated on familiar objects in those spaces. This pairing of personal querying history with recurring usage locations turns routine contexts into a lightweight information provider.
@@ 4 Preliminary Evaluation
To assess the practicality and potential of proactive, context-aware AR assistance, we conducted a two-part preliminary evaluation: (1) a technical analysis of Memento capabilities and performance, and (2) a “take-home” diary-based user study examining user acceptance, user experience, real-world usage patterns, and how the technology integrates into everyday routines.
@@ 4.1 Technical Evaluation
Identifying users’ action contexts, and delivering timely and accurate situated AR content, are crucial for the practical adoption of proactive context-aware systems. In this section, we assess the context-aware capabilities of Memento and report our results.
Spatial activity context matching accuracy. We evaluate the accuracy of the Spatial Activity Recognizer in identifying spatial and activity contexts from egocentric views using text-based scene representations. Each image is converted into a textual description with GPT-4o, and CLIP is used to generate embeddings directly from these textual summaries. To assess classification performance, we compute centroid embeddings for each spatial and activity class using a sampled dataset of 1,000 egocentric images drawn from Ego4D. For each test sample, its embedding is compared against class centroids using cosine similarity, retrieving the closest match as the predicted label.
The recognizer achieves an F1 score of 0.82 for Scene context classification and 0.82 for Activity context classification (see Table 1). We find that this tendency continues until images contain motion blur or low lighting, where the accuracy drops to an F1 score of 0.57 and 0.30, respectively. These results indicate that text-based embeddings can provide semantic distinction to identify recurring spatial activities for proactive resurfacing, even without fine-tuning or visual modality fusion.
Interaction referent recognition accuracy. We found that the open-vocabulary object detector, YOLO-World-XL, occasionally struggles to robustly localize a referent that the user is interacting with. While it outperforms DetCLIP-T in speed (52 vs 2FPS) with similar accuracy (34.4 vs 35.4 AP), false positives led to faulty retrieval of RSAM and inconsistent situating of proactive AR visualizations, impacting the user experience.
@@
Table 1: Performance comparison of CLIP variants–CNN-based (ResNet-101) and Transformer-based (ViT-B/32, ViT-L/14) architectures, for spatial Scene and Activity recognition tasks.
@@ 4.2 User Evaluation and Feedback
We collected data from 25 sessions conducted over a 2-4 day period, resulting in a total of $\approx$435 aggregated minutes, and 196 memories (the total number of query-generated memories across sessions; including user-dismissed memories) across all participants. Although each session lasted for 17.4 minutes on average, and was relatively short in duration, participants were exposed to 135 memories via proactive recalls (68.9% of total memories).
@@
Participants and procedure. We recruited 9 (P1-P9) self-volunteered participants (7 male, 2 female), aged 25-36 ($\mu$=29.1, $\sigma$=3.7), from Computer Science, Mathematics, and Physics backgrounds. Prior AR familiarity was broadly distributed on a 5-point scale (1:Not familiar, 5:Familiar), ($\mu$=2.8, $\sigma$=1.3), and gesture-based interaction experience showed a similar spread ($\mu$=3.1, $\sigma$=1.6). All participants reported regular daily use of personal computing devices. Participants’ identities were anonymized, and each provided informed consent prior to participation. This study was conducted under the IRB approval of Stony Brook University ($1173920$). The participants were instructed to complete a brief pre-study questionnaire and a 30-minute introduction to Memento. They were told to use Memento naturally over 2-4 days in their daily indoor environments, with each usage counted as a “session.” After each session, participants submitted short diary entries describing context, perceived usefulness, and ease of integration, including a 5-point Likert scale and open-ended feedback. System logs of spatial context, interactions, and resurfaced memories were collected in parallel. After the multi-day use, participants completed a post-study survey comparing proactive and reactive support. Semi-structured interviews followed to explore how proactive cues affected routines and what limitations they encountered. To ensure consistency across participants, proactive resurfacing sensitivity was fixed at one resurfacing per 24 hours.
@@
Repeated daily usages. Participants praised Memento for its proactive task reminding ability–“I liked that I could place reminders around the house. Next to my shoes, I’ll know I need to go out for a run”(P3). However, one user reported referent misalignment: “It mistook a square grocery bag to be a luggage and suggested me travel time”(P3). A post-study interview revealed that the user moved close to the referent, enforcing our recognition model to infer with minimal visual context. In another case, a spatial activity was over-generalized: “It said I’m gazing at an object. It’s too generic”(P2). We discuss candidate solutions in Section 5.
@@
Acceptability in daily routine. Users were favorable towards Memento becoming part of their daily routine: “Each time I use it, I get more comfortable with the interaction style. It feels like a solid foundation”(P5). However, the requests to extend Memento, to a lightweight glasses-based AR were suggested across participants (P1, P3, P4, P6, P8); “It’ll be nice to have a VR app integrated into lifestyle through glasses that pops up information quick”(P6). A non-tech user said : “It is cool, but I wear a huge glasses, so I can’t wear it for long”(P3). Furthermore, even with a multi-day AR experience, users expressed the need for a longer duration, to observe any changes in their routines: “I cannot say it changed my routines yet. It clearly helped me, but I’d need to try it on for longer. I might feel differently if it is a glasses I can wear casually, like the Ray-ban”(P1); “Very promising system. Looking forward to seeing how it adapts with more use”(P5). We extend our discussion to Section 5.
@@
Proactive vs Reactive assistance. Proactive context-aware AR in Memento was evaluated positively: “I check weather and traffic every morning, this gave me the answers before I checked”(P7); “It reduces a step in your daily patterns”(P6); “Great for reminders. I usually use Google Calendar, but many times my phone is on silent or put away, I’m unable to follow through. Proactivity definitely helps more and put bunch of To-Dos around my house”(P3). Another participant compared Memento to a traditional reactive assistant: “I use Google Nest Mini to voice query and get responses, but Memento recall of previous queries is very useful”(P1), and praised on its simplicity “Biggest pros is that I don’t need to think about setting up the conditions, just automatic and simple”(P6).
5 Future Work and Discussion
@@
Mixed-initiative context control and activity granularity. Accurate resurfacing depends on inferring when a resurfaced memory is contextually appropriate. AR content that is mis-triggered may negate the usefulness and utility of proactive visualization. While participants generally appreciated autonomous triggers, they wanted occasional corrective control: “I sometimes wanted more control”(P5). This suggests that proactive AR should adopt mixed-initiative mechanisms that remain simple, yet adjustable. Prior work on user-assistant collaboration and conditional triggers, such as IFTTT (If-This-Then-That) [ifttt, shakeri2023user]-like structures, may provide contexts for lightweight refinements. Improving recognition accuracy may also benefit from hierarchical activity representations as suggested by ongoing efforts in structured activity inference [edge2024local, userllm, xie2024embodied, yang2025egolife], mitigating over-generalization and helping maintain alignment between resurfacing timing and user intent.
Continuity of context across realities. Participants responded positively when resurfaced information appeared, “where it felt natural,” especially when memories were tied to objects already associated with an activity. While the current panel-based overlays provide invariance to layout changes, they lack deeper geometric coupling with the environment. Closer integration between AR content and spatial referents could strengthen relevance and improve interpretability and explainability, building on design considerations for spatially anchored AR interfaces [lee2023design]. Future work should explore referent-aware embeddings that balances spatial robustness with meaningful anchoring.
Need for longitudinal routine pattern evaluation. Our diary-based deployment captured short-term routine patterns but could not reveal how resurfacing should adapt as habits evolve. Participants occasionally treated the system as a temporary “study task,” leading to usage that did not fully reflect everyday rhythms. Longer-term field studies would allow users to internalize proactive AR as part of daily life, similar to long-term wearable research [bressa2022data]. Understanding when proactive reminders should scale back, pause, or revive requires extended observation of routine drift and habit change over weeks or months, rather than days.
@@ 6 Conclusion
To explore the usability of proactive, context-aware assistance on a wearable AR device, we designed a foundational system, Memento, that captures and stores users’ verbal interactions as Referent-anchored Spatiotemporal Activity Memories, and proactively resurfaces them based on recurring daily contexts. We introduced a hybrid data retrieval structure combining spatial indexing and semantic vector search, and implemented a simplified context-aware system by integrating multimodal sensing, LLM-based reasoning, and open-vocabulary referent detection. Through real-world use cases, we investigated the potential for everyday use of a proactive AR assistant, and discuss takeaways and future work.